Previous

Calendar

August

1999

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

1

JD213
Site Visit (AT): Details

at 6:30 MST, data system back up

rain guage precip data start being more than daily avg (ie, valid 30-min data start HERE).

added NR.LITE net radiometer (pointed toward SE on 2nd level from top of tower)

2

JD214

3

JD215
Site Visit (AT): Details

replaced N2 cylinder

Changed profiler (li-6251) dessicant

swiped krypton hygrometer

on 8/2, RH at 2m was not working correctly (removed sensor, dried it out and resealed it). Seemed ok after repairs.

4

JD216

5

JD217

6

JD218

7

JD219

8

JD220

9

JD221

10

JD222
Site Visit (AT): Details

N2 cylinder was empty (must of had a leak)

at 16:00, new span tank?

from 10:00-11:00, re-installed the repaired LI-6262 (sn #638) in the EC system

Autocalibrations have not been working (since either 8/3 or 8/6)

at 12:00, autocals running again

needed to power-cycle LI-6251 (in prof system)

11

JD223

12

JD224
Site Visit (AT): Details

Profiler has been off for ~24 hrs

at 10:00, put IRGA #368 back into profiler system and running again

Autocalibrations still having problem (delivery pressure on regulator too low?)

13

JD225

14

JD226

15

JD227

16

JD228

17

JD229

18

JD230

19

JD231
Site Visit (AT): Details

from 10:22-11:14, manual calibration of LI-6251/prof system

at 11:33, changed filter in EC system (also switched out teflon tubing in LI-6262 box for dekabon tubing; no change in flow rate)

at 12:00, change prep.config and cover.config to include O3 fast sensors

at 14:30, took down REBS #2 (#1 still up)

in PM, top circuit breaker popped

20

JD232

21

JD233

22

JD234
Site Visit (AT): Details

from 11:15-16:00, working on fast O3 system

at 16:00, changed nupro inlet filter for EC system

23

JD235

24

JD236

25

JD237
Site Visit (AT): Details

Re-did plumbing on SECLOD O3 system

Only using fast1.o3 (disconnected pump to other)

In CNR/Blanken CR23X (start using CNR2.CSI)

26

JD238

27

JD239

28

JD240

29

JD241

30

JD242
Site Visit (AT): Details

at top of tower, replaced the BLANKEN CR23X with the BOWLING CR23X (still called CNR/RAD CR23X logger)

re-leveled radiation sensors: CNR-1, Q7.1 (this was okay), PAR sensors, NR.Lite (also moved, pointed toward south 1 level below CNR-1 and Q7.1)

31

JD243
Site Visit (LSD): Details

Laura moved the 1m co2 inlet to 11cm (0.11m), also moved it out into the woods

1

JD244

2

JD245

3

JD246

4

JD247